OpenStudy (clamin):

PLEASE HELP!! MEDAL!! Consider the function f(x)=3x^2+3. What is the average rate of change of f(x) over the interval [5,9]? a. 168 b. 52 c. 42 d. 21

OpenStudy (anonymous):

it might be c but I'm not sure :/

OpenStudy (clamin):

ohh okay, i wanna know how you got the answer..please

geerky42 (geerky42):

\(average~rate~of~change ~from~a~to~b= \dfrac{f(b)-f(a)}{b-a}\)

geerky42 (geerky42):

So here you have \(\dfrac{f(9)-f(5)}{9-5}\)

OpenStudy (clamin):

how do i solve this?? @geerky42

geerky42 (geerky42):

\[f(\color{red}x)=3\color{red}x^2+3\]So \(f(\color{red}9)=3(\color{red}9)^2+3\) and \(f(\color{red}5)=3(\color{red}5)^2+3\)